Plane with the wind: 240 Miles in 2 hours against the wind:240 miles in 3 hours find rate of plane and wind.
Answer by checkley77(12844)

D=RT
240=(R-W)3 AGAINST THE WIND.
R-W=240/3
R-W=80
240=(R+W)2 WITH THE WIND.
R+W=240/2
R+W=120
R-W=80
(120-80)/2=40/2=20 MPH FOR THE WIND SPEED.
PROOF:
240=80*3
240=240
240=120*2
240=240